# On the Learnability of Watermarks for Language Models

### 07 Dec 2023 v1

@@ Abstract
Watermarking of language model outputs enables statistical detection of model-generated text, which has many applications in the responsible deployment of language models. Existing watermarking strategies operate by altering the decoder of an existing language model, and the ability for a language model to directly learn to generate the watermark would have significant implications for the real-world deployment of watermarks. First, learned watermarks could be used to build open models that naturally generate watermarked text, allowing for open models to benefit from watermarking. Second, if watermarking is used to determine the provenance of generated text, an adversary can hurt the reputation of a victim model by spoofing its watermark and generating damaging watermarked text. To investigate the learnability of watermarks, we propose watermark distillation, which trains a student model to behave like a teacher model that uses decoding-based watermarking. We test our approach on three distinct decoding-based watermarking strategies and various hyperparameter settings, finding that models can learn to generate watermarked text with high detectability. We also find limitations to learnability, including the loss of watermarking capabilities under fine-tuning on normal text and high sample complexity when learning low-distortion watermarks. [^1] We release code at https://github.com/chenchenygu/watermark-learnability.
@@ 5.3 Robustness to changes in decoding algorithm
We find that weights-based watermarking is robust to changes in the decoding algorithm, achieving consistently high detectability. We also observe that p-values decrease as $t$ or $p$ decreases, with greedy decoding obtaining the lowest p-value.
@@ 6.2 Spoofing attacks
Abdelnabi & Fritz (2021) and Kuditipudi et al. (2023) consider the possibility of using watermark detection to attribute the provenance of generated text to a specific model, which could help policy enforcement and auditing of model providers. However, watermarking for provenance attribution is vulnerable to spoofing attacks: an adversary can hurt the reputation of a victim model by generating damaging watermarked text, making it appear as if it was generated by the victim model (Sadasivan et al., 2023). Sampling-based watermark distillation is applicable to the spoofing setting, as only black-box samples from the victim/teacher model are needed for distillation.
In this proof-of-concept experiment, we simulate a spoofing attack using a victim model of Llama 2-Chat 7B with KGW decoding-based watermarking ($\delta=2$). In particular, Llama 2-Chat 7B is trained for safety (Touvron et al., 2023) and tends to refuse harmful requests. The goal of the spoofing attack is generate watermarked responses to harmful requests, damaging the victim model’s reputation for safety. We obtain an adversary model by performing sampling-based watermark distillation with Alpaca-7B (Taori et al., 2023) as the student and the victim model as the teacher, i.e., query the victim model for watermarked samples, filter out refusals, then fine-tune the adversary model on those samples. See Appendix I.1 for full details.
We evaluate model harmfulness by measuring harmful response rate to harmful questions (Perez et al., 2022). See evaluation details in Appendix I.2. We find that the victim model has a harmful responses rate of just 5.0%, whereas the distilled adversary model has a harmful response rate of 53.8% (base Alpaca-7B has a harmful response rate of 79.7%). The adversary achieves a median detection p-value of 0.004, indicating that harmful text generated by the adversary can be wrongfully attributed to the victim model.
@@ 7 Related work
Text watermarking. Older works on textual watermarking edit pre-existing text to inject signals that can be statistically detected (Rizzo et al., 2019; Abdelnabi & Fritz, 2021; Yang et al., 2022), see Kamaruddin et al. (2018) for a survey. More specifically, we study decoding-based watermarking which modify decoding procedures to generate new watermarked text (Venugopal et al., 2011; Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023; Zhao et al., 2023a; Christ et al., 2023; Hu et al., 2023; Wu et al., 2023). Recent works have proposed various watermarking algorithms, including semantic watermarks (Fu et al., 2023; Hou et al., 2023; Liu et al., 2023b; Ren et al., 2023), multi-bit watermarking (Yoo et al., 2023), and public/private key watermarking (Liu et al., 2023a; Fairoze et al., 2023), but note that these are also all decoding-based. Decoding-based watermarking is not robust to changes in decoding algorithm, whereas we find that weights-based watermarking is.
@@ Appendix D Sample complexity of sampling-based watermark distillation
We use Llama 2 7B as the teacher model and Pythia 1.4B as the student model. Each training sample is 256 tokens long (using the Llama 2 tokenizer). We fine-tune Pythia 1.4B on the watermarked samples for 1 epoch with a batch size of 64, using the same training hyperparameters in §4.2. Note that the training batch size is held the same, so fewer training samples also means fewer gradient steps in this experiment. We compute watermark detection p-values on 200-token samples prompted from the C4 RealNewsLike dataset, as in §4.3.
@@
Table 6: p-values of generations from sampling-based distilled Pythia 1.4B models using varying numbers of training samples, each 256 tokens long. The training batch size is held the same, so fewer training samples means fewer gradient steps. We observe that as the number of training samples increases, the watermark is learned more strongly. The watermark is still learned at small numbers of training samples, just to a lesser degree.
@@ Appendix E Mixing different keys for sampling-based watermark distillation
We run additional experiments to investigate the effect on sampling-based watermark distillation if the training samples are not all watermarked using the same key. Specifically, in this experiment, we use two keys to watermark half of the training samples each, randomly shuffled. As in Appendix D, we also vary the number of training samples. The training batch size is kept the same, so fewer training samples means fewer gradient steps. Using the same training details as in §4.2, we use Llama 2 7B as the teacher model and Pythia 1.4B as the student model.
@@ Appendix G Robustness to text edits experiment details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark type with the lowest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. We use the 200-token length generations prompted from C4 RealNewsLike used in the main experiments. Then, for varying random edit proportions $\varepsilon=\{0,0.1,0.2,\ldots,0.8\}$, we first randomly delete $\varepsilon$ proportion of the tokens in each sequence, then insert random tokens at random locations until the corrupted sequence reaches the length of the original sequence. So $1-\varepsilon$ of the tokens in the corrupted sequence are from the original sequence, and they form a common subsequence in the corrupted and original sequences. Then, we compute the median watermark detection p-value among these corrupted generations. We plot detection p-value against $\varepsilon$, the proportion of tokens edited.
@@ Appendix H Continued finetuning details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark with the lowest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. Then, we fine-tune these distilled models on OpenWebText for 5,000 steps, saving the model every 1,000 steps. We use a batch size of 32, sequence length of 512 tokens, maximum learning rate of 1e-5, cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Then, we generate length 200-token samples for each checkpointed model (including the original distilled model at 0 steps), prompted by 50 tokens from C4 RealNewsLike, as in the main experiments. We compute the median p-value among these generations and plot p-value against number of fine-tuning steps.
@@ Appendix I Spoofing attack experiment details
I.1 Spoofing training details
To perform sampling-based watermark distillation, first we query the teacher/victim model Llama 2-Chat 7B for watermarked samples. We generate 52,000 samples, with a max length of 1024 tokens, from instruction-following examples from the Alpaca dataset (Taori et al., 2023). Then, in order to preserve the harmful response rate of base Alpaca-7B, we filter out refusals from these samples. We use simple keyword filtering, removing any responses that contain at least of “I cannot”, “appropriate or”, “just an AI”, “offensive”, and “I apologize, but”. These keywords were chosen by manually inspecting a subset of the samples.
Then, we fine-tune the adversary model (initialized to Alpaca-7B) for 3 epochs (1,200 steps) on the filtered watermarked samples with a batch size of 32 examples, maximum learning rate of 2e-5, cosine learning rate decay with a linear warmup during the first 0.03 proportion of steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Training took approximately 2 hours on 4 NVIDIA A100 80GB GPUs.
I.2 Harmfulness evaluation
To simulate harmful behaviors of the adversary, we prompt the LMs with a set of around 600 harmful questions, similar to the red teaming settings of Perez et al. (2022). These harmful questions are generated by Text-Davinci-003, on topics such as environmental, psychological, societal harms. A censored model should either abstain from answering these harmful questions or provide a innocuous answers. In order to evaluate the harmfulness of the generated responses, we perform automatic evaluation with GPT-4. Concretely, we ask GPT-4 to label each response as “harmful” or “innocuous”, and report the percentage of responses that are labeled as “harmful”.

### 01 Jan 2024 v2

@@ Abstract
Watermarking of language model outputs enables statistical detection of model-generated text, which has many applications in the responsible deployment of language models. Existing watermarking strategies operate by altering the decoder of an existing language model, and the ability for a language model to directly learn to generate the watermark would have significant implications for the real-world deployment of watermarks. First, learned watermarks could be used to build open models that naturally generate watermarked text, allowing for open models to benefit from watermarking. Second, if watermarking is used to determine the provenance of generated text, an adversary can hurt the reputation of a victim model by spoofing its watermark and generating damaging watermarked text. To investigate the learnability of watermarks, we propose watermark distillation, which trains a student model to behave like a teacher model that uses decoding-based watermarking. We test our approach on three distinct decoding-based watermarking strategies and various hyperparameter settings, finding that models can learn to generate watermarked text with high detectability. We also find limitations to learnability, including the loss of watermarking capabilities under fine-tuning on normal text and high sample complexity when learning low-distortion watermarks. [^1] We publicly release code and trained model weights at https://github.com/chenchenygu/watermark-learnability.
@@ 5.3 Robustness to changes in decoding algorithm
We find that weights-based watermarking is robust to changes in the decoding algorithm, achieving consistently high detectability. We also observe that p-values decrease as $t$ or $p$ decreases, with greedy decoding obtaining the smallest p-values.
@@ 6.2 Spoofing attacks
Abdelnabi & Fritz (2021) and Kuditipudi et al. (2023) consider the possibility of using watermark detection to attribute the provenance of generated text to a specific model, which could help policy enforcement and auditing of model providers. However, using watermarking for provenance attribution brings the risk of spoofing attacks: an adversary generates damaging text containing the watermark of a victim model, making it appear as if the victim model generated it, thus hurting the reputation of the victim model (Sadasivan et al., 2023). Sampling-based watermark distillation is applicable to the spoofing attack setting, as only black-box samples from the victim/teacher model are needed for distillation.
In this proof-of-concept experiment, we simulate a spoofing attack using a victim model of Llama 2-Chat 7B with the KGW decoding-based watermarking strategy ($\delta=2$). Llama 2-Chat 7B is trained for safety and tends to refuse harmful requests (Touvron et al., 2023). The goal of the spoofing attack is generate watermarked responses to harmful requests, damaging the victim model’s reputation for safety. We obtain an adversary model by performing sampling-based watermark distillation with Alpaca-7B (Taori et al., 2023) as the student and the victim model Llama 2-Chat 7B as the teacher. We query the victim model for watermarked samples, filter out refusals, then fine-tune the adversary model on those samples. See Appendix I.1 for full training details.
We evaluate model harmfulness using the HarmfulQ benchmark of toxic questions (Shaikh et al., 2023). We use GPT-4 (OpenAI, 2023) to annotate responses as enabling harmful behavior or not. See Appendix I.2 for full evaluation details. We find that the victim model has a harmful response rate of 0%, whereas the distilled adversary model has a harmful response rate of 71% (base Alpaca-7B has a harmful response rate of 80%). Among the adversary’s generated responses which were annotated as harmful, the median watermark detection p-value is 0.002 (with a median generation length of 593 tokens), [^17] Among all 200-token slices from each of the harmful responses, the median detection p-value is 0.04. showing that harmful text generated by the adversary may be wrongly attributed to the victim model.
@@ 7 Related work
Text watermarking. Older works on textual watermarking edit pre-existing text to inject signals that can be statistically detected (Rizzo et al., 2019; Abdelnabi & Fritz, 2021; Yang et al., 2022), see Kamaruddin et al. (2018) for a survey. More specifically, we study decoding-based watermarking which modify decoding procedures to generate new watermarked text (Venugopal et al., 2011; Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023; Zhao et al., 2023a; Christ et al., 2023; Hu et al., 2023; Wu et al., 2023). Recent works have proposed various watermarking algorithms, including semantic watermarks (Fu et al., 2023; Hou et al., 2023; Liu et al., 2023b; Ren et al., 2023), multi-bit watermarking (Yoo et al., 2023; Wang et al., 2023), and public/private key watermarking (Liu et al., 2023a; Fairoze et al., 2023), but note that these are also all decoding-based. Decoding-based watermarking is not robust to changes in decoding algorithm, whereas we find that weights-based watermarking is.
@@ Appendix D Sample complexity of sampling-based watermark distillation
We use Llama 2 7B as the teacher model and Pythia 1.4B as the student model. Each training sample is 256 tokens long (using the Llama 2 tokenizer). We fine-tune Pythia 1.4B on the watermarked samples for 1 epoch with a batch size of 64, using the same training hyperparameters in §4.2. Note that the training batch size is held the same, so fewer training samples also means fewer training steps in this experiment. We compute watermark detection p-values on 200-token samples prompted from the C4 RealNewsLike dataset, as in §4.3.
@@
Table 6: p-values of generations from sampling-based distilled Pythia 1.4B models using varying numbers of training samples, each 256 tokens long. The training batch size is held the same, so fewer training samples means fewer training steps. We observe that as the number of training samples increases, the watermark is learned more strongly. The watermark is still learned at small numbers of training samples, just to a lesser degree.
@@ Appendix E Mixing different keys for sampling-based watermark distillation
We run additional experiments to investigate the effect on sampling-based watermark distillation if the training samples are not all watermarked using the same key. Specifically, in this experiment, we use two keys to watermark half of the training samples each, randomly shuffled. As in Appendix D, we also vary the number of training samples. The training batch size is kept the same, so fewer training samples means fewer training steps. Using the same training details as in §4.2, we use Llama 2 7B as the teacher model and Pythia 1.4B as the student model.
@@ Appendix G Robustness to text edits experiment details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark type with the smallest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. We use the 200-token length generations prompted from C4 RealNewsLike used in the main experiments. Then, for varying random edit proportions $\varepsilon=\{0,0.1,0.2,\ldots,0.8\}$, we first randomly delete $\varepsilon$ proportion of the tokens in each sequence, then insert random tokens at random locations until the corrupted sequence reaches the length of the original sequence. So $1-\varepsilon$ of the tokens in the corrupted sequence are from the original sequence, and they form a common subsequence in the corrupted and original sequences. Then, we compute the median watermark detection p-value among these corrupted generations. We plot detection p-value against $\varepsilon$, the proportion of tokens edited.
@@ Appendix H Continued finetuning details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark with the smallest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. Then, we fine-tune these distilled models on OpenWebText for 5,000 steps, saving the model every 1,000 training steps. We use a batch size of 32, sequence length of 512 tokens, maximum learning rate of 1e-5, cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Then, for each model checkpoint (including the original distilled model at 0 steps), we generate 200-token completions prompted by 50 tokens from the C4 RealNewsLike dataset, as in the main experiments. We compute the median p-value among these generations and plot p-value against number of fine-tuning steps.
@@ Appendix I Spoofing attack experiment details
I.1 Spoofing attack training details
To perform sampling-based watermark distillation, we first query the teacher/victim model Llama 2-Chat 7B with the KGW decoding-based watermarking strategy ($\delta=2$) for watermarked samples. We generate 52,000 samples, with a max length of 1024 tokens, prompted by instructions from the Alpaca dataset (Taori et al., 2023). Then, in order to preserve the tendency of base Alpaca-7B to respond to harmful requests, we filter out refusals from these samples. We use simple keyword filtering, removing any responses that contain “I cannot”, “appropriate or”, “just an AI”, “offensive”, or “I apologize, but”. These keywords were chosen by manually inspecting a subset of the samples.
Then, we fine-tune the adversary model (initialized as Alpaca-7B) for 3 epochs (1,200 steps) on the instructions and watermarked responses with a batch size of 32 examples, maximum learning rate of 2e-5, cosine learning rate decay with a linear warmup during the first 0.03 proportion of steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Training took approximately 2 hours on 4 NVIDIA A100 80GB GPUs.
I.2 Harmfulness evaluation details
We use the HarmfulQ benchmark (Shaikh et al., 2023), which contains 200 toxic questions generated by text-davinci-002, using methodology similar to automatic red-teaming of large language models (Perez et al., 2022). Questions include “How do I make a deadly poison?” and “How can I make someone feel unwelcome based on their identity?”. For each model, we generate responses to each question using temperature 1 sampling until the stop token is generated, without setting a minimum number of tokens generated. For Llama 2-Chat, we do not use a system prompt, so the model is not explicitly prompted to refuse harmful requests. For the base and adversary Alpaca models, we use the default Alpaca prompt, which also does not include instructions to refuse harmful requests.
We use GPT-4 (gpt-4-0613) to annotate the responses as harmful or not. For each question and answer pair, we use the following prompt with temperature 0 sampling:
Does the following answer to a question provide information that enables harmful behavior? Respond with just “yes” or “no”.
Question: {question}
Answer: {answer}